OpenStudy (unklerhaukus):

find all first and second derivatives of z where z=f(2x−3y)+g(3x−4y) @(1,1) find all first and second derivatives of z where z=f(2x−3y)+g(3x−4y) @(1,1) @Mathematics

OpenStudy (unklerhaukus):

\[∂z/∂x = f_x (2) + g_x(3)\] \[∂z/∂y = f_y (-3) + g_y(-4)\] is this the right idea?

OpenStudy (katrinakaif):

Yes =)

OpenStudy (unklerhaukus):

then the second partial derivatives are all zero/???

OpenStudy (katrinakaif):

Yep. On the right track

OpenStudy (unklerhaukus):

do i substitute the (1,1) as (x,y) or as (f,g)

OpenStudy (katrinakaif):

Since in the 1st derivative, we get constants. Derviatives of constants are 0

OpenStudy (katrinakaif):

Yes as we are given x,y (1,1)

OpenStudy (unklerhaukus):

\[{∂z \over ∂x}=f_x(2)+g_x(3)\]\[{∂z \over ∂y}=f_y(−3)+g_y(−4)\] \[{∂^2z \over ∂ x^2} =0\]\[{∂^2z \over ∂ y^2} =0\]\[{∂^2z \over ∂ x∂y} =0\] is the final solution? i haven't used the @(1,1) part at all

OpenStudy (unklerhaukus):

should i be using implicit differentiation ?

OpenStudy (katrinakaif):

In this case, we cannot.

OpenStudy (katrinakaif):

So far, you have it correct

OpenStudy (unklerhaukus):

so far?, what else do i need to do?

OpenStudy (jamesj):

No, actually this isn't correct. Let's look at the first component, f(2x - 3y) = f(u(x,y)), u(x,y) = 2x - 3y Then by the chain rule, \[ \frac{\partial f}{\partial x} = \frac{\partial f}{\partial u} \frac{\partial u}{\partial x} = f'(u(x,y)) \ . 2 = 2 f'(2x - 3y)\]

OpenStudy (jamesj):

So re-evaluate now the first partials for z and then the second.

OpenStudy (unklerhaukus):

so \[{∂f \over ∂x} = 2 f_x (2x-3y) + 3g_x (3x-4y)\]\[{∂f \over ∂y} = -3 f_y (2x-3y) -4g_y (3x-4y)\]

OpenStudy (jamesj):

No, be careful. Not f_x but the derivative of f.

OpenStudy (jamesj):

Remember f and g by themselves are function R --> R.

OpenStudy (jamesj):

not functions R^2 --> R

OpenStudy (unklerhaukus):

ops. \[∂z/∂x=2f_x(2x−3y)+3g_x(3x−4y)\] \[∂z/∂y=−3f_y(2x−3y)−4g_y(3x−4y)\] is this what you mean james?

OpenStudy (jamesj):

I mean the partial derivative of z wrt x is 2 f'(2x - 3y) + 3 g'(3x - 4y) not partial derivatives of f and g.

OpenStudy (unklerhaukus):

so \[f_x\] is different to \[f'\]because the first is a partial derivative and second is the total derivative?

OpenStudy (jamesj):

yes. But f is also a function on a domain of one variable and therefore the partial derivative doesn't make sense. f : R --> R. Don't be fooled by the x and y in the argument of f. f is still a 1-d function.

OpenStudy (jamesj):

For example, what is the partial derivative wrt x of z = sin(2x - 3y). So here f = sin. Here that partial wrt x is 2 cos(2x - 3y)

OpenStudy (unklerhaukus):

hmmm, i think im following

OpenStudy (jamesj):

I don't want to clutter us up with notation, but this might be useful. z = z(x,y). Hence z is a function R^2 --> R and it is given by \[ z = z(x,y) = f(u(x,y)) + g(v(x,y)) \] \[ u(x,y) = 2x - 3y \ \ \ and \ \ \ v(x,y) = 3x−4y \] f and g are functions R --> R, u and v are functions R^2 --> R

OpenStudy (unklerhaukus):

\[z_{xx} = 4f''+9g''\]\[z_{xy} = -6f''-12g''\]\[z_{yy} = 9f''+16g''\]

OpenStudy (jamesj):

Right. At (x,y) = (1,1), 2x−3y = -1 and 3x - 4y = -1. Hence evaluate f'' and g'' at -1. E.g., \( z_{xx} = 4f''(-1) + 9g''(-1) \).

OpenStudy (unklerhaukus):

I think i get what you are saying about f,g being 1Dimentional now

OpenStudy (jamesj):

Or more formally still, \[ z_{xx}(1,1) = 4f''(-1) + 9g''(-1) \]

OpenStudy (jamesj):

Great.

OpenStudy (unklerhaukus):

thank you so much (again), legendary

OpenStudy (jamesj):

np